Hi
I tried this but can't find the solution.
Solve for x,y by substitution.
8x-y=13
y=-8x+1

Answer by Alan3354(69443)

Solve for x,y by substitution.
8x-y=13
y=-8x+1
----------------
Sub for y in the 1st eqn
8x -(-8x+1) = 13
16x - 1 = 13
x = 7/8
----
y = -6